# Runge-Kutta Approximations for Direct Coning Compensation Applying Lie Theory

### v1

@@ 5 Results
To quantify solver accuracy a reference trajectory is required. We define $\boldsymbol{\omega}(t)$ analytically such that it may be evaluated exactly at infinite resolution in $t$. Given $\boldsymbol{\omega}(t)$, and initial condition $\boldsymbol{\theta}(t_{k})$, the attitude trajectory is obtained by numerically solving the ODE in eq. (13). No effort was made to define a trajectory matching a particular vehicle. Instead, our goal is to demonstrate our ability to recover the solution $\boldsymbol{\phi}(t)$ from discrete samples of $\boldsymbol{\omega}$ or $\Delta\boldsymbol{\theta}$ rather than the infinite resolution available to our reference solution.
Applying sequential rotations using the convention from eq. (31), we know that
@@
We seek to evaluate $T\left(\Delta\boldsymbol{\phi}_{k}\right)$ by solving eq. (13) from an initial condition $\boldsymbol{\phi}\left(t_{k}\right)=0$. Solving this ODE numerically using gyro measurements we first consider direct access to instantaneous rate measurements $\boldsymbol{\omega}\left(t_{k}\right),\boldsymbol{\omega}\left(t_{k}+\Delta t/2\right),\boldsymbol{\omega}\left(t_{k}+\Delta t\right)$. Instantaneous rates are natural inputs to the generalized solver Algorithm 1. In contrast we also consider approximating $\boldsymbol{\omega}$ from measurements $\Delta\boldsymbol{\theta}_{k-1},\Delta\boldsymbol{\theta}_{k},,\Delta\boldsymbol{\theta}_{k+1}$ provided by integrating rate gyros.
Results are depicted in Figure 3. In the legend, $\omega$ and $\Theta$ are used to distinguish results obtained from instantaneous rates or integrated rate measurements, respectively. The subscript on $\Theta$ indicates the number of integrated measurements used in modeling $\boldsymbol{\omega}$. Specifically, $\Theta_{3}$ requires $\Delta\boldsymbol{\theta}_{k+1}$ such that $\Delta\boldsymbol{\phi}_{k}$ is a lagged update. Recall that the SingleSpeed $\Theta_{2}$ algorithm form eq. (51) is identical to the classical single-speed algorithm from eq. (30).
Figure 3: Algorithm performance on discrete measurements. ExMid, RK3 and RK4 are employ the Runge-Kutta algorithm 1 with Tableaux 2, 3, and 4, respectively. SingleSpeed $\Theta_{2}$ (traditional coning compensation) and $\Theta_{3}$ use Eqs. (51) and (55), respectively.
@@
Results for FwdElr, ExMid, and RK4 operating on direct samples of $\boldsymbol{\omega}$ illustrate expected reduction in solver error function with solver order. The improvement in solver error from $\Theta_{2}$ to $\Theta_{3}$ illustrates the impact of a quadratic model for $\boldsymbol{\omega}$. A quadratic model is required to unlock the full benefits of a 4th order model as indicated by the agreement between RK4 $\boldsymbol{\omega}$ and SingleSpeed $\Theta_{3}$. Without the "future" measurement $\Delta\boldsymbol{\theta}_{k=1}$, we observe SingleSpeed $\Theta_{2}$ approximation performs similarly to RK3 $\boldsymbol{\omega}$. Further, the agreement suggests negligible impact of the approximation eq. (16) for our example trajectory and justifies the simplification eq. (30).
@@ 6 Conclusion
This work introduces a new framework for developing coning correction algorithms of varying order and with a varying number of gyro measurements. The framework is developed from a Lie-theoretic perspective and uses rigorous approximations of Runge-Kutta methods to solve the requisite differential equations. The classical single-speed coning algorithm from eq. (30) popularized by Miller in the early 1980s is found to be a special case of an RK4 integrator. To fully realize the benefits of a 4th order model, however, requires the addition of an additional prior gyro measurement. The methods presented in this work may be straightforwardly extended to higher-order (using the appropriate Butcher tableau) and to more measurements.
@@ Acknowledgments
This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government. SAND2025-13567O
Appendix A Analytic Derivation of Single-Speed Coning Correction
@@
over the interval $t\in[-\Delta t,\Delta t]$ and where $\{\emph{{p}}_{i}\}_{i=1}^{2}\in\mathbb{R}^{3}$ are constant vectors that we do not know. Now, we may use (58) to analytically compute $\boldsymbol{\theta}(t)$ from eq. (20) as
@@
where the constant of integration is $\emph{{c}}=0$ since our initial condition is $\boldsymbol{\theta}(0)=\textbf{0}$. It follows, therefore, that
@@
where we have made use of the cross-product identity $\emph{{p}}_{2}\times\emph{{p}}_{1}=-\emph{{p}}_{1}\times\emph{{p}}_{2}$. It is now straightforward to analytically compute the coning integral from eq. (27)
@@
The trouble is that we do not know $\emph{{p}}_{1}$ and $\emph{{p}}_{2}$ and, therefore, cannot (yet) evaluate eq. (61) to find $\Delta\boldsymbol{\beta}$. To circumvent this problem, first compute the actual gyro measurement at the current and prior time steps as
@@
from which we observe that
@@
We recognize the final relationship to be precisely the unknown term in the coning integral from eq. (61). Therefore, substituting this result we compute the coning correction to be

### v2

@@ 5 Results
Applying sequential rotations using the convention from eq. (32),
@@
Here, $T\left(\Delta\boldsymbol{\phi}_{k}\right)\in SO(3)$ is parameterized by $\Delta\boldsymbol{\phi}$ which is defined in eq. (20). The algorithms presented above may be used to compute the parameters $\Delta\boldsymbol{\phi}$ by solving the ODE of eq. (13) from an initial condition $\boldsymbol{\phi}\left(t_{k}\right)=0$. Reconstructions $\Delta\hat{\boldsymbol{\phi}}$ are obtained from discrete samples of $\boldsymbol{\omega}$ or $\Delta\boldsymbol{\theta}$. Algorithm performance is scored against a truth trajectory, $\Delta\boldsymbol{\phi}^{*}$. The complete simulation is available on GitHub [HITMAN_2026].
Instead of defining $\Delta\boldsymbol{\phi}^{*}$ analytically, Bézier curves were used to define $\boldsymbol{\omega}(t)$ analytically. One benefit is that $\Delta\boldsymbol{\theta}$, given by eq. (22), are available in closed form. Two Bézier curves were defined for $t\in\left[0,1\right]$: one benign and one challenging. A relatively flat cubic polynomial was chosen for the benign curve. For the challenging curve, control points were selected by randomly generating a $6\times 3$ matrix with standard Gaussian values. For each curve, segments were selected with a random (uniformly sampled) starting location and varying step sizes. Discrete uniformly sampled measurements about the segement of interest were collected in $\Omega$ and $\Theta$ for instantaneous rate and integrated rate measurements, respectively.
Truth trajectories were computed by solving the ODE eq. (13) numerically to obtain $\Delta\boldsymbol{\phi}^{*}$. Specifically, SciPy solve_ivp was used with the default RK4(5) variable step-size solver, tolerances reduced to 1e-10, and a max step size of 1e-4. Variable step-size solvers sample $\boldsymbol{\omega}(t)$ with fine resolution in contrast to the coarse measurements available to algorithms under test. Alternatively, $\boldsymbol{\phi}(t)$ could be defined analytically in which case integrated measurements $\Delta\boldsymbol{\theta}$ must be evaluated numerically. No appreciable changes were observed in contrasting algorithm performance when choosing to define $\boldsymbol{\phi}(t)$ analytically instead of $\boldsymbol{\omega}(t)$.
For instantaneous rate measurements, attitude updates were computed by solving the ODE (33) numerically with algorithm 1. Named algorithms FwdElr, ExMid, RK3 and RK4 employed the Runge-Kutta algorithm 1 with Tableaux 3, 4, 5, and 6, respectively. For each tableau, the first column indicates the required sample points of $\boldsymbol{\omega}$. As a matter of notation, the subscript decoration on $\Omega_{n}\in\mathbb{R}^{3\times n}$ is used to partition only the first $n$ columns of $\Omega$ defined in eq. (47). Notice FwdElr and ExMid operate on $\Omega_{1}$ and $\Omega_{2}$, respectively. Both RK3 and RK4 operate on $\Omega_{3}$.
For integrated rate measuremments, attitude updates are computed using single-speed algorithms. The SingleSpeed $\Theta_{2}$ algorithm from eq. (52) is identical to the classical single-speed algorithm from eq. (31). It operates on $\Theta_{2}$ defined in eq. (45). The SingleSpeed $\Theta_{3}$ uses eq. (56), operating on $\Theta_{3}$ defined in eq. (53).
Direct comparison of $\Delta\hat{\boldsymbol{\phi}}$ are $\Delta\boldsymbol{\phi}^{*}$ are problematic due to the non-uniqueness of exponential coordinates. Mapping exponential coordinates back to their group elements avoids artifacts due rotations of $2\pi$. When quantifying algorithm performance, the Frobenius norm of the group elements was employed
$\left\|\exp\left(\left(\Delta\hat{\boldsymbol{\phi}}\right)^{\wedge}\right)-\exp\left(\left(\Delta\boldsymbol{\phi}^{*}\right)^{\wedge}\right)\right\|_{\textrm{Fro}}$ (59)
@@
Results are depicted in Figure 7. In the legend, $\Omega$ and $\Theta$ are used to distinguish results obtained from instantaneous rates or integrated rate measurements, respectively. The subscripts indicate the number of measurements utilized. Results for FwdElr, ExMid, RK3, and RK4 operating on direct samples of $\boldsymbol{\omega}$ illustrate expected reduction in solver error function with solver order. The improvement in solver error from $\Theta_{2}$ to $\Theta_{3}$ illustrates the impact of a quadratic model for $\boldsymbol{\omega}$. A quadratic model for $\boldsymbol{\omega}$ is required to recover benefits of a 4th order integrator as indicated by the agreement between RK4 $\boldsymbol{\omega}$ and SingleSpeed $\Theta_{3}$ for the benign trajectory. Without the "future" measurement $\Delta\boldsymbol{\theta}_{k=1}$, the SingleSpeed $\Theta_{2}$ approximation performs similarly to RK3 $\boldsymbol{\omega}$. Further, the agreement suggests negligible impact of the approximation eq. (16) for the example trajectory and justifies the simplification eq. (31) even on challenging trajectories.
(a) Benign Trajectory
(b) Challenging Trajectory
Figure 7: Algorithm performance on discrete measurements for benign and challenging trajectories. Attitude propagation from instantaneous or integrated rate measurements are indicated by $\Omega_{n}$ and $\Theta_{n}$, respectively where the subscript $n$ indicates the number of measurements utilized. SingleSpeed $\Theta_{2}$ (traditional coning compensation) and $\Theta_{3}$ use eqs. (52) and (56), respectively.
@@ 6 Conclusion
This work introduces a new framework for developing coning correction algorithms of varying order and with a varying number of gyro measurements. The framework is developed from Lie theory and approximations of Runge-Kutta methods to solve the requisite differential equations. The classical single-speed coning algorithm popularized by Miller in the early 1980s is found to be a special case of an RK4 integrator. To fully realize the benefits of a 4th order model, however, requires the addition of an additional prior gyro measurement. The methods presented in this work may be straightforwardly extended to higher-order (using the appropriate Butcher tableau) and to more measurements.
@@ Acknowledgments
This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.
Appendix: Analytic Derivation of Single-Speed Coning Correction
@@
over the interval $t\in[-\Delta t,\Delta t]$ and where $\{\emph{{p}}_{i}\}_{i=1}^{2}\in\mathbb{R}^{3}$ are constant vectors that are unknown. Now, (60) may be used to analytically compute $\boldsymbol{\theta}(t)$ from eq. (21) as
@@
where the constant of integration is $\emph{{c}}=0$ since the initial condition is $\boldsymbol{\theta}(0)=\textbf{0}$. It follows, therefore, that
@@
which made use of the cross-product identity $\emph{{p}}_{2}\times\emph{{p}}_{1}=-\emph{{p}}_{1}\times\emph{{p}}_{2}$. It is now straightforward to analytically compute the coning integral from eq. (28)
@@
The trouble is that $\emph{{p}}_{1}$ and $\emph{{p}}_{2}$ are unknown and, therefore, it is not (yet) possible to evaluate eq. (63) to find $\Delta\boldsymbol{\beta}$. To circumvent this problem, first compute the actual gyro measurement at the current and prior time steps as
@@
from which it is observed that
@@
The final relationship is recognized to be precisely the unknown term in the coning integral from eq. (63). Therefore, substituting this result to compute the coning correction effects